### v1

# Density control of multi-agent swarms via bio-inspired leader-follower plasticity

Abstract
The design of control systems for the spatial self-organization of mobile agents is an open challenge across several engineering domains, including swarm robotics and synthetic biology. Here, we propose a bio-inspired leader-follower solution, which is aware of energy constraints of mobile agents and is apt to deal with large swarms. Akin to many natural systems, control objectives are formulated for the entire collective, and leaders and followers are allowed to plastically switch their role in time. We frame a density control problem, modeling the agents’ population via a system of nonlinear partial differential equations. This approach allows for a compact description that inherently avoids the curse of dimensionality and improves analytical tractability. We derive analytical guarantees for the existence of desired steady-state solutions and their local stability for one-dimensional and higher-dimensional problems. We numerically validate our control methodology, offering support to the effectiveness, robustness, and versatility of our proposed bio-inspired control strategy.
keywords:
Collective behavior; density control; leader-follower.
thanks: Corresponding authors Mario di Bernardo and Maurizio Porfiri.
1 Introduction
Across a wide range of applications, from swarm robotics [1] to collective construction [2], environmental management [3], synthetic biology [4], and search-and-rescue operations [5], engineers are to design the spatial displacement of large swarms. Leader-follower strategies [6] are commonly used toward this objective, usually assuming that a population of controller agents (leaders) is given the task of inducing a desired behavior in another population (followers). The implementation of these leader-follower strategies is often informed by our understanding of the mechanisms underpinning collective behavior of various animal species, such as fish [7, 8], birds [9], and humans [10, 11]. Despite considerable success, existing engineered systems fail to capture the richness of their biological counterparts in two key aspects.
First, leaders and followers are part of a fixed hierarchy, which does not capture role switching due to behavioral plasticity. Behavioral plasticity is the reversible biological mechanism behind changes in the behavior of organisms due to internal or environmental stimuli [12]. For instance, in fish schools behavioral plasticity arises when followers become leaders as they access novel information about predators and food locations [13, 14]. Likewise, in flocks of migrating birds, it occurs as leaders step down from frontal positions in the formation to recover energy [15]. Similar observations are gathered in human groups performing collective tasks, such as games [16] and team sports [17].
Second, leaders in engineered systems are typically viewed as fixed reference signals in open-loop, not receptive to the unfolding collective dynamics [18]. Even when a feedback loop is included in the design of the leader-follower control strategy, its goal is only to steer the behavior of followers without taking into account the entire collective [19]. In biological systems, the behavior of leaders adapts over time in response to collective needs; for example, foraging or predator avoidance are key to survival and reproduction of both leaders and followers [20].
Addressing both of these gaps would benefit several application areas related to spatial organization of large swarms. For example, a recent perspective [21] has noted that the implementation of plasticity in robotic swarms may be a viable option to face the high unpredictability of the real world and grant agents the necessary flexibility needed to perform a wide spectrum of tasks. Enabling spontaneous roles’ switching between leaders and followers may be conducive to an overall reduction of control efforts during task execution, thereby containing energy costs for the swarm. Such a benefit may further extend to the domain of synthetic biology, where there is a need for control schemes that could regulate the composition of microbial consortia composed to guarantee efficient labor division [22]. Here, we develop a continuum model for a large population of behaviorally plastic leaders and followers agents solving a density control problem.
Mathematically, the control of large swarms introduces several challenges: ($i$) the state space grows exponentially with the number of agents; ($ii$) communication graphs may be time-varying and constrained; and ($iii$) individual agent dynamics are coupled through nonlocal interaction terms. Continuum approaches using partial differential equations (PDEs) offer a promising avenue to circumvent these difficulties by modeling agent densities rather than individual states, thus achieving dimension reduction while preserving essential collective dynamics [23, 24, 25]. Mean-field models of interacting populations of mobile agents have gained substantial momentum over the last decade [26, 27, 28], as they offer compact formulations to increase computational and analytical tractability. Bio-inspired switching mechanisms between leaders and followers have been explored in [29, 30, 31, 32, 33] for modeling purposes, but none of these efforts has presented feedback control actions to induce desired collective behaviors. A mean-field optimal control problem with transient leadership has been formulated in [34], but, due to the absence of closed form feedback solutions, it can only be numerically approximated, yielding limited insight that may inform control design in realistic settings. Related work on leader selection [35, 36] and switching mechanisms [37] exists, but these approaches either maintain fixed population assignments or focus on synchronization rather than spatial density control.
We draw insight from the literature on reacting mixtures [38], used to describe blood flows [39] and tumors growth [40]. In particular, we model plasticity as a chemical reaction taking place between two fluids, associated with the continuum description of leaders and followers. We assume that plasticity is not common to all agents, where some of them may not be allowed switching their role. Whether or not an agent is a leader or a follower is not distinguishable by the rest of the group. We derive conditions for the existence and stability of the solution, in terms of key model parameters that can be part of engineering design, such as the fraction of non-plastic agents, interaction kernels, and desired densities. Our approach enables us to provide explicit feedback control laws that guarantee exponential convergence, stability analysis for the coupled PDE system, and feasibility conditions through adaptive leader selection.
The main control-theoretic contributions of the paper are as follows: ($i$) We derive explicit feedback control laws for density regulation in heterogeneous swarms with adaptive leader-follower role assignment, providing exponential convergence guarantees. We include plastic leaders and followers, which can dynamically exchange roles, and non-plastic followers, which cannot switch to leadership roles; ($ii$) we establish necessary and sufficient conditions for the existence and local stability of desired equilibrium density distributions under role-switching dynamics; ($iii$) we extensively validate our strategy through numerical experiments, demonstrating that adaptive role switching enhances system robustness to parametric uncertainties compared to fixed-hierarchy approaches.
The rest of the paper is organized as follows. In Sec. 2, we present the problem statement for a one-dimensional (1D) scenario. The proposed control strategy is formulated in Sec. 3 and its numerical validation is detailed in Sec. 4. An agent-based model that supports the use of our continuum approach is presented in Sec. 5. The extension to higher dimensions and the corresponding numerical validation are expounded in Sec. 6. Section 7 concludes the manuscript summarizing our main findings and proposing avenues of future research.
2 Problem Statement
Our control objective is to steer the spatial distribution of the entire collective toward a desired configuration through coordinated leader actions, while allowing adaptive role assignment within the swarm. We adopt a continuum approach to describe the densities of the agents. The mathematical formulation consists of three coupled convection-diffusion equations on the unit circle $\mathcal{S}=[-\pi,\pi]$,
$\rho^{L}_{t}(x,t)+\left[\rho^{L}(x,t)u(x,t)\right]_{x}+$ $\left[\rho^{L}(x,t)(f*\rho)(x,t)\right]_{x}$
$=D$ $\rho^{L}_{xx}(x,t)+q(x,t),$ (1a)
$\rho^{F}_{t}(x,t)+\big{[}\rho^{F}(x,t)(f*\rho)(x,t)$ $\big{]}_{x}=D\rho^{F}_{xx}(x,t)$
$\qquad\quad-q(x,t),$ (1b)
$\eta^{F}_{t}(x,t)+\big{[}\eta^{F}(x,t)(f*\rho)(x,t)$ $\big{]}_{x}=D\eta^{F}_{xx}(x,t).$ (1c)
Here, $(\cdot)_{t}$ and $(\cdot)_{x}$ indicate partial derivatives with respect to time and spatial coordinates, respectively. These equations model the spatio-temporal dynamics of the density of three subsets of agents, respectively
  • leaders (whose density is $\rho^{L}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that ($i$) react to the rest of the group, ($ii$) are controlled through the periodic velocity field $u:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$, and ($iii$) may become followers due to the reacting mechanism $q:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$;
  • plastic followers (whose density is $\rho^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that ($i$) react to the rest of the population, and ($ii$) may become leaders due to the reacting mechanism $q$;
  • non-plastic followers (whose density is $\eta^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that only react to the rest of the population but cannot switch their role to leaders.
We consider all-to-all inter- and intra-population interactions taking place through a periodic interaction kernel $f$ (soft-core and vanishing with distance). We model these interactions through a cross-convectional non-local term involving $\rho=\rho^{L}+\rho^{F}+\eta^{F}$. To account for the presence of noise affecting the behavior of any agent in the real world, we include a diffusion process, weighted by the diffusion coefficient $D$.
By summing (1a), (1b) and (1c), we obtain the dynamics of the density of the whole population
$\rho_{t}(x,t)+\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}=\\ =D\rho_{xx}(x,t),$ (2)
which is independent of the reacting term $q$. By imposing periodic boundary conditions for (1a), (1b) and (1c),
$\rho^{i}(-\pi,t)$ $=\rho^{i}(\pi,t),\;\forall t\in\mathbb{R}_{\geq 0},\,i=L,F$ (3a)
$\eta^{F}(-\pi,t)$ $=\eta^{F}(\pi,t),\;\forall t\in\mathbb{R}_{\geq 0},$ (3b)
we ensure that the total population mass is conserved. In particular, we have
$\left(\int_{\mathcal{S}}\rho(x,t)\,\mathrm{d}x\right)_{t}=-\left[\rho^{L}(x,t)u(x,t)\right]_{-\pi}^{\pi}\\ -\left[\rho(x,t)(f*\rho)(x,t)\right]_{-\pi}^{\pi}+D\left[\rho_{x}(x,t)\right]_{-\pi}^{\pi}=0,$ (4)
because of periodicity. Notice that (1c) satisfies mass conservation, but (1a) and (1b) do not, due to the reacting term $q$. If $\int_{\mathcal{S}}q\,\mathrm{d}x$ equals 0, there is no net mass transfer between leaders and followers.
Without loss of generality, we normalize the total mass to one by setting
$\int_{\mathcal{S}}\rho(x,t)\,\mathrm{d}x=M^{L}(t)+M^{F}(t)+\Phi^{F}=1,$ (5)
where $M^{L}$ is the leaders’ mass, $M^{F}$ is the followers’ mass, and $\Phi^{F}$ is the constant mass of non-plastic followers. We define the fraction of the population that is allowed to switch role, $p$, as
$p=1-\Phi^{F}.$ (6)
Equation set (1) is complemented by the following initial conditions:
$\rho^{i}(x,0)$ $=\rho^{i}_{0}(x),\,i=L,F$ (7a)
$\eta^{F}(x,0)$ $=\eta^{F}_{0}(x).$ (7b)
We consider the density control problem of choosing $p$, $u$, and $q$ in (1) so that the population density asymptotically converges almost everywhere (a.e.) towards a desired time-invariant density profile $\bar{\rho}:\mathcal{S}\to\mathbb{R}_{\geq 0}$,
$\lim_{t\to\infty}\|\bar{\rho}(\cdot)-\rho(\cdot,t)\|_{2}=0,$ (8)
where $\|\cdot\|_{2}$ denotes the $\mathcal{L}^{2}$ norm over $\mathcal{S}$. We note that the control problem pertains to the density of the entire collective, comprising leaders and followers. As an additional control specification capturing energy costs, we tune the steady-state leaders-to-followers mass ratio to a desired value $\hat{r}$, namely,
$\lim_{t\to\infty}\frac{M^{L}(t)}{M^{F}(t)}=\hat{r}.$ (9)
3 Bio-inspired Control
3.1 Design of the leaders’ velocity field
We define the error function for the density of the entire collective,
$e(x,t)=\bar{\rho}(x)-\rho(x,t).$ (10)
Using (2), the error dynamics reads
$e_{t}(x,t)=\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}\\ -D\rho_{xx}(x,t),$ (11)
with periodic boundary conditions and initial conditions that can be derived from (3) and (7).
Theorem 1 (Global exponential convergence).
Assume $\rho^{L}>0$ for any $x\in\mathcal{S}$ and $t\in\mathbb{R}_{\geq 0}$. Choosing the control input $u$ in (1a) according to
$\left[\rho^{L}(x,t)u(x,t)\right]_{x}=-Ke(x,t)-\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}\\ +D\rho_{xx}(x,t),$ (12)
where $K>0$ is a control gain, the error converges globally and exponentially to 0 point-wise in $\mathcal{S}$, that is,
$e(x,t)=e(x,0)\,\mathrm{exp}(-Kt).$ (13)
{pf}
Substituting (12) into (11), we obtain
$e_{t}(x,t)=-Ke(x,t),$ (14)
which is linear and does not involve spatial derivatives. Its analytical solution yields (13).∎
Remark 2.
The control input $u$ can be found by spatial integration of (12), as
$u(x,t)=\frac{1}{\rho^{L}(x,t)}\bigg{[}-K\int e(x,t)\,\mathrm{d}x$
$-\rho(x,t)(f*\rho)(x,t)$ $+D\rho_{x}(x)\bigg{]},$ (15)
which is well-defined only for $\rho^{L}>0$. Note that $u$ can be shown to be periodic (see Corollary 1 in [19]).
Remark 3.
Under the hypothesis of Theorem 1 and using (13), we establish the following closed-form expression for the density of the collective:
$\rho(x,t)=\bar{\rho}(x)\left[1-\mathrm{exp}(-Kt)\right]+\rho_{0}(x)\mathrm{exp}(-Kt),$ (16)
where $\rho_{0}=\rho^{L}_{0}+\rho^{F}_{0}+\eta^{F}_{0}$.
3.2 Design of the reacting term
Our design of the reacting term $q$ in (1a) and (1b) is driven by two main objectives: ($i$) to ensure that the hypothesis about the strict positivity of $\rho^{L}$ in Theorem 1 holds and ($ii$) to achieve a desired leaders-to-followers’ mass ratio at steady state (9).
We achieve these goals by choosing the reacting term $q$ as
$q(x,t)=\frac{1}{2}\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\frac{1}{2}\left[\rho^{*}(x,t)(f*\rho)(x,t)\right]_{x}\\ -\frac{D}{2}\rho^{*}_{xx}(x,t)+g(x,t),$ (17)
where
$\rho^{*}:=\rho^{L}-\rho^{F},$ (18)
and $g$ obeys to the mass action law
$g(x,t)=K_{FL}\rho^{F}(x,t)-K_{LF}\rho^{L}(x,t).$ (19)
The positive reaction rates $K_{FL}$ and $K_{LF}$ represent the propensity of leaders to become followers and of followers to become leaders, respectively.
Theorem 4 (Strict positivity of $\rho^{L}$).
With $u$ chosen as in (12) and $q$ as in (17), $\bar{\rho}$ is a steady-state solution of (1) with $\bar{\rho}^{L},\bar{\rho}^{F}>0$ and $\bar{\eta}^{F}\geq 0$ for any $x\in\mathcal{S}$, if and only if
$p>\hat{p}=1-\min_{x}\left[\bar{\rho}(x)\frac{\int_{\mathcal{S}}h(x)\,\mathrm{d}x}{h(x)}\right],$ (20)
with
$h(x)=\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(x)\,\mathrm{d}x\right].$ (21)
{pf}
[…]
$\rho^{*}_{t}(x,t)=2q(x,t)-\left[\rho^{L}(x,t)u(x,t)\right]_{x}\\ -\left[\rho^{*}(x,t)(f*\rho)(x,t)\right]_{x}+D\rho^{*}_{xx}(x,t).$ (22)
$\rho^{L}$ and $\rho^{F}$ can be recovered from $\rho$, $\rho^{*}$, and $\eta^{F}$ through the change of variables
$\rho^{L}(x,t)$ $=\frac{1}{2}\left[\rho(x,t)+\rho^{*}(x,t)-\eta^{F}(x,t)\right],$ (23a)
$\rho^{F}(x,t)$ $=\frac{1}{2}\left[\rho(x,t)-\rho^{*}(x,t)-\eta^{F}(x,t)\right].$ (23b)
Substituting (17) into (22), and using (23) yields
$\rho^{*}_{t}(x,t)=-a\,\rho^{*}(x,t)+b\,\rho(x,t)-b\,\eta^{F}(x,t),$ (24)
where
$a$ $:=K_{FL}+K_{LF},$ (25a)
$b$ $:=K_{FL}-K_{LF}.$ (25b)
Under the control action discussed in Theorem 1, $\bar{\rho}$ is a steady-state solution for (2), so that we look for steady-state solutions of (24) and (1c). We start by considering (1c) with $\eta^{F}_{t}=0$, $\eta^{F}(x,t)=\bar{\eta}^{F}(x)$, and $\rho(x,t)=\bar{\rho}(x)$, which gives
$D\bar{\eta}^{F}_{xx}(x)-\left[\bar{\eta}^{F}(x)(f*\bar{\rho})(x)\right]_{x}=0.$ (26)
Integrating (26) twice in space (see Appendix A) yields
$\bar{\eta}^{F}(x)=\frac{\Phi^{F}}{\int_{\mathcal{S}}h(x)\,\mathrm{d}x}h(x).$ (27)
We remark that $\bar{\eta}^{F}$ is positive, periodic, and $\int_{\mathcal{S}}\bar{\eta}^{F}\,\mathrm{d}x=\Phi^{F}$ by construction (see Appendix A for more details). We can now find the steady-state of $\rho^{*}$ by setting $\rho^{*}_{t}=0$, $\rho(x,t)=\bar{\rho}(x)$, $\rho^{*}(x,t)=\bar{\rho}^{*}(x)$, $\eta^{F}(x,t)=\bar{\eta}^{F}(x)$ in (24). This gives
$\bar{\rho}^{*}(x)=\frac{b}{a}\left[\bar{\rho}(x)-\bar{\eta}^{F}(x)\right].$ (28)
Hence, using (23), at steady-state we obtain
$\bar{\rho}^{L}(x)=\frac{1}{2}\left[\bar{\rho}(x)\left(1+\frac{b}{a}\right)-\bar{\eta}^{F}(x)\left(1+\frac{b}{a}\right)\right],$ (29a)
$\bar{\rho}^{F}(x)=\frac{1}{2}\left[\bar{\rho}(x)\left(1-\frac{b}{a}\right)-\bar{\eta}^{F}(x)\left(1-\frac{b}{a}\right)\right].$ (29b)
Since $|b/a|<1$ by construction, $\bar{\rho}^{L}$ and $\bar{\rho}^{F}$ are strictly positive if
$\bar{\rho}(x)>\bar{\eta}^{F}(x),\;\forall\,x\in\mathcal{S},$ (30)
which is satisfied under condition (20) (substituting (27) into (30), and recalling (6)).
($\implies$) The existence of a steady-state solution for (1) with $\rho^{L},\rho^{F}>0$ and $\bar{\eta}^{F}$ implies that
$\bar{\rho}^{L}(x)+\bar{\rho}^{F}(x)>0,\;\forall\,x\in\mathcal{S}.$ (31)
By adding and subtracting $\bar{\eta}^{F}$, we obtain
$\bar{\rho}^{L}(x)+\bar{\rho}^{F}(x)+\bar{\eta}^{F}(x)>\bar{\eta}^{F}(x),\;\forall\,x\in\mathcal{S},$ (32)
which is equivalent to
$\bar{\rho}(x)>\bar{\eta}^{F}(x),\;\forall\,x\in\mathcal{S}.$ (33)
Substituting (27) into (33), and recalling (6), completes the proof. ∎
Remark 5.
Theorem 4 gives conditions about the minimum fraction of agents that can switch role such that $\bar{\rho}$ can be a meaningful steady-state solution for (1), that is (5) and (8) hold with $\rho^{L}>0$, $\rho^{F},\eta^{F}\geq 0$. Hence, it consists in an existence result for the solutions of the problem given in Sec. 2.
Corollary 6.
The requirement in (9) can be ensured by choosing $K_{LF}$ and $K_{FL}$ in (19).
{pf}
From (29), we compute the steady-state leaders-to-followers mass ratio, that is
$\frac{\int_{\mathcal{S}}\bar{\rho}^{L}(x)\,\mathrm{d}x}{\int_{\mathcal{S}}\bar{\rho}^{F}(x)\,\mathrm{d}x}=\frac{K_{FL}}{K_{LF}}.$ (34)
Hence, by appropriately choosing the reacting rates $K_{LF}$ and $K_{FL}$, we fulfill (9).∎
Example 7.
We provide an illustration of how to use Theorem 4. We consider a population of agents interacting via the periodic Morse interaction kernel (long-range attraction, short-range repulsion)
$f(x)=\frac{1}{L_{r}}f_{r}(x)-\frac{\alpha}{L_{a}}f_{a}(x)$ (35)
where $L_{a}$ and $L_{r}$ are the length scales of the attractive and repulsive part of the interaction kernel, and
$f_{i}(x)=\frac{\mathrm{sgn(x)}}{\mathrm{exp}\left(\frac{2\pi}{L_{i}}\right)-1}\left[\mathrm{exp}\left(\frac{2\pi-|x|}{L_{i}}\right)-\mathrm{exp}\left(\frac{|x|}{L_{i}}\right)\right],$ (36)
with $i=L,R$ (see [41] for more details). We set $L_{a}=\pi$, $\alpha=2$, $L_{r}=\pi/6$, and $D=0.05$. We select the desired density for the group to be a von Mises, that is,
$\bar{\rho}(x)=Z\,\mathrm{exp}\left[k\cos(x-\mu)\right],$ (37)
where we fix mean $\mu=0$ and concentration coefficient $k=1$. $Z$ is chosen so that $\bar{\rho}$ integrates to 1. Using (20), we establish that the fraction of agents that are allowed to switch the role should be larger than $\hat{p}\approx 0.15$ for the desired collective density to be feasible.
(a) $t=0$
(b) $t=t_{\mathrm{f}}$
Figure 1: Bimodal regulation. (a,b) Initial/final (solid black) and desired (dashed black) density of the collective. In the inset, we report the initial/final densities of leaders (solid blue), plastic followers (solid orange), and non-plastic followers (solid purple) along with the density predictions at steady state from Theorem 4 for the three populations (dashed and same color coding). (c) Time evolution of the KL divergence (top panel) and leaders’ and followers’ mass (bottom panel). (d) Final distribution profile of the leaders’ velocity $u$ (top panel) and reacting term $q$ (bottom panel).
3.3 Stability analysis
In this section, we assess the stability properties of our control solution, exerted through the velocity field $u$ in (2) and the reactive term $q$ in (17). From Theorem 1, we know that, if $\rho^{L}>0$, global convergence of $\rho$ toward $\bar{\rho}$ is ensured. Theorem 4 instead gives conditions for $\rho^{L}$ to be striclty positive at steady-state. Hence, we now prove local stability of the solution whose existence is proved in Theorem 4. Let us recall the function $e=\bar{\rho}-\rho$, and define error functions
$e^{*}(x,t)$ $:=\bar{\rho}^{*}(x)-\rho^{*}(x,t),$ (38a)
$e^{\eta}(x,t)$ $:=\bar{\eta}^{F}(x)-\eta^{F}(x,t),$ (38b)
with $\bar{\rho}^{*}$ defined in (28) and $\bar{\eta}^{F}$ in (27).
Theorem 8 (Local stability).
Under the conditions of Theorem 4, error functions (10), (38a), and (38b) locally converge to 0 almost everywhere if
$\|\bar{\rho}_{x}(\cdot)\|_{2}<\frac{2D}{\|f(\cdot)\|_{2}}$ (39)
{pf}
The error dynamics under the effect of $u$ and $q$ are given by
$e_{t}(x,t)=-K\,e(x,t),\qquad\qquad$ (40a)
$e^{*}_{t}(x,t)=-a\,e^{*}(x,t)+b\,e(x,t)$ $-b\,e^{\eta}(x,t),$ (40b)
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)(f*\bar{\rho})(x)]_{x}-$ $\left[e^{\eta}(x,t)(f*e)(x,t)\right]_{x}=$
$=De^{\eta}_{xx}(x,t)-[\bar{\eta}^{F}(x)$ $(f*e)(x,t)]_{x}.$ (40c)
The first two equations of the error system are linear and do not involve spatial derivatives. The third equation, however, is nonlinear and involves spatial derivatives.
By linearizing the last equation about the origin, we find
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)(f*\bar{\rho})(x)]_{x}=$
$=De^{\eta}_{xx}(x,t)-[\bar{\eta}^{F}(x)$ $(f*e)(x,t)]_{x}.$ (41)
We substitute (13) into (3.3), yielding
$e^{\eta}_{t}(x,t)+[e^{\eta}(x,t)$ $(f*\bar{\rho})(x)]_{x}=De^{\eta}_{xx}(x,t)$
$+\mathrm{exp}(-Kt)\left[\bar{\eta}^{F}(x)v^{0}(x)\right]_{x},$ (42)
where $v^{0}=f*e^{0}$ with $e^{0}(x)=e(x,0)$.
We introduce the Lyapunov functional
$V(t)=\|e^{\eta}(\cdot,t)\|_{2}^{2}.$ (43)
The time derivative of $V$ can be expressed as
$V_{t}(t)=2\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{t}(x,t)\,\mathrm{d}x=\\ =2D\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{xx}(x,t)\,\mathrm{d}x\\ -2\int_{\mathcal{S}}e^{\eta}(x,t)\left[e^{\eta}(x,t)(f*\bar{\rho})(x)\right]_{x}\,\mathrm{d}x\\ -2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x,$ (44)
where we substituted (3.3) and $\tilde{v}=(\bar{\eta}^{F}v^{0})_{x}$ We expand the fist term on the right-hand side of (44) as
$2D\int_{\mathcal{S}}e^{\eta}(x,t)e^{\eta}_{xx}(x,t)\,\mathrm{d}x=-2D\int_{\mathcal{S}}\left(e^{\eta}_{x}(x,t)\right)^{2}\,\mathrm{d}x=\\ =-2D\|e_{x}^{\eta}(\cdot,t)\|_{2}^{2},$ (45)
where we applied integration by parts (recalling the periodicity of the functions). We similarly expand the second term on the right-hand side of (44) as
$-2\int_{\mathcal{S}}e^{\eta}(x,t)\left[e^{\eta}(x,t)(f*\bar{\rho})(x)\right]_{x}\,\mathrm{d}x=\\ =2\int_{\mathcal{S}}e^{\eta}_{x}(x,t)e^{\eta}(x,t)(f*\bar{\rho})(x)\,\mathrm{d}x=\\ =\int_{\mathcal{S}}\left[\left(e^{\eta}(x,t)\right)^{2}\right]_{x}(f*\bar{\rho})(x)\,\mathrm{d}x=\\ =-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x,$ (46)
where we used integration by parts (twice), and exploited the identity $\left[\left(e^{\eta}\right)^{2}\right]_{x}=2e^{\eta}e^{\eta}_{x}$. Substituting (45) and (46) into (44), we obtain
$V_{t}(t)=-2D\|e_{x}^{\eta}$ $(\cdot,t)\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x$
$-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x.$ (47)
By using Poincaré-Wirtinger inequality (see Lemma 2 in [19]), we can bound this as
$V_{t}(t)\leq-2D\|e^{\eta}$ $(\cdot,t)\|_{2}^{2}-\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x$
$-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x.$ (48)
For the second term on the right-hand side of (3.3), we have
$\left|\int_{\mathcal{S}}\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\,\mathrm{d}x\right|\leq\\ \leq\int_{\mathcal{S}}\left|\left(e^{\eta}(x,t)\right)^{2}(f*\bar{\rho})_{x}(x)\right|\,\mathrm{d}x=\\ =\|e^{\eta}(\cdot,t)e^{\eta}(\cdot,t)(f*\bar{\rho})_{x}(\cdot)\|_{1}\leq\\ \leq\|e^{\eta}(\cdot,t)\|_{2}\|e^{\eta}(\cdot,t)\|_{2}\|(f*\bar{\rho})_{x}(\cdot)\|_{\infty}\leq\\ \leq\|e^{\eta}(\cdot,t)\|_{2}^{2}\|f(\cdot)\|_{2}\|\bar{\rho}_{x}(\cdot)\|_{2},$ (49)
[…]
$\left|-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}(x,t)\tilde{v}(x)\mathrm{d}x\right|\leq\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}(\cdot,t)\tilde{v}(\cdot)\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}(\cdot,t)\|_{2}\|\tilde{v}(\cdot)\|_{2}$ (50)
By applying bounds (49) and (50) to (3.3), we obtain
$V_{t}(t)\leq\left(-2D+\|f(\cdot)\|_{2}\|\bar{\rho}_{x}(\cdot)\|_{2}\right)V(t)\\ +2\mathrm{exp}(-Kt)\|\tilde{v}(\cdot)\|_{1}\sqrt{V(t)}.$ (51)
If $\|\bar{\rho}\|_{2}<2D/\|f\|_{2}$, the right-hand side in (51) converges to 0 thanks to Lemma 4 in [19] (with $\beta=-2D+\|f\|_{2}\|\bar{\rho}_{x}\|_{2}$ $\gamma=0$, and $\delta=2\|\tilde{v}\|_{2}$ ). Hence, by Comparison Lemma [42], we know $\|e^{\eta}\|_{2}^{2}$ (locally) converges to 0.
Since (40a) converges point-wise and (40c) converges locally in $\mathcal{L}^{2}(\mathcal{S})$, we can analyze (40b), rewriting it as
$e^{*}_{t}(x,t)=-ae^{*}(x,t)+w(x,t),$ (52)
where $w$ is a bounded function converging to 0 asymptotically in time and a.e. in $\mathcal{S}$. Computing the unilateral Laplace transform in time to (52) yields
$E^{*}(x,s)=\frac{W(x,s)}{s+a}.$ (53)
where $E^{*}$ and $W$ are Laplace transform of $e^{*}$ and $w$, respectively. Given that $a>0$, the application of the final value theorem yields
$\lim_{t\to\infty}e^{*}(x,t)=\lim_{s\to 0}\frac{sW(x,s)}{s+a}=0,$ (54)
where we used the fact that $\lim_{s\to 0}sW(x,s)=0$ since $w$ asymptotically converges to zero. ∎
4 Numerical Validation
We validate our theoretical results through numerical simulations, demonstrating the effectiveness of our control strategy. We characterize performance through the time evolution of the Kullback-Leibler (KL) divergence between $\rho$ and $\bar{\rho}$, that is,
$\mathcal{D}_{KL}(t)=\int_{\mathcal{S}}\rho(x,t)\log\left(\frac{\rho(x,t)}{\bar{\rho}(x)}\right)\,\mathrm{d}x.$ (55)
As a steady-state performance index, we consider $\mathcal{D}_{KL}^{ss}=\mathcal{D}_{KL}(t_{\mathrm{f}})$, where $t_{\mathrm{f}}$ is the final instant of the simulation.
4.1 Bimodal regulation
Figure 2: Robustness analysis to perturbations in (a) diffusion coefficients and (b) parameters of the interaction kernels. For different values of $p$, we show $\mathcal{D}_{KL}^{ss}$ (blue) and leaders’ mass (orange) at steady-state (in solid gray the predicted minimum plasticity ensuring feasibility).
[…]
4.2 Robustness analysis
Next, we demonstrate how the fraction of plastic agents affects robustness with respect to perturbations. We consider a desired von Mises distribution ($\kappa=1$, $\mu=0$), Morse interactions as in (35)($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=2$), $K=10$, and $D=0.02$. We introduce perturbations to either the diffusion coefficient – doubling it for the followers in Eqs. (1c) and (1b) with respect to its nominal value used for the leaders in (1a)– or the interaction kernel parameters – reducing $L_{a}$ by 20% and increasing $L_{r}$ by 20% with respect to their nominal values for all the agents. Starting from equilibrium configurations [^1] For values of $p$ below the minimum threshold prescribed by Theorem 4, steady-state configurations $\bar{\rho}^{L}$ and $\bar{\rho}^{F}$ are negative in some regions of the domain. For these cases, we translate initial configurations upwards to become non-negative and re-normalize them to a predefined mass., we assess performance degradation for different values of $p$. When the perturbation affects only followers (as for the test with respect to $D$), we choose $K_{LF}$ and $K_{FL}$ to ensure the steady-state leaders’ mass is constant across different values of $p$. This makes the amplitude of the perturbation constant when varying $p$. Results in Figs. 2(a) and 2(b) show that above the minimum threshold for $p$ predicted in Theorem 4, agents rearrange to counteract perturbations, maintaining steady-state performance. Below this threshold, performance degrades significantly.
5 Agent-based Model
To bridge the gap between our continuum theoretical framework and practical implementation, we develop a discrete agent-based model that captures the essential PDE dynamics at the individual agent level. This validates the continuum approximation and demonstrates implementability in realistic multi-agent systems.
The agent-based model consists of coupled stochastic differential equations describing agent positions, with stochasticity from Gaussian noise and role-switching according to reaction rates.
In particular, we choose
$\mathrm{d}x_{i}(t)=\Bigg{[}\frac{1}{N^{LF}}$ $\sum_{i=1}^{N^{LF}}f(\{x_{i}(t),x_{j}(t)\})$
$+\frac{1}{M}\sum_{i=1}^{M}$ $f(\{x_{i}(t),y_{j}(t)\})+u_{i}(t)\lambda_{i}(t)\Bigg{]}\mathrm{d}t$
$+\sqrt{2D}\,\mathrm{d}W_{i},\;i=1,\dots,N^{LF},$ (56a)
$\mathrm{d}y_{i}(t)=\Bigg{[}\frac{1}{M}\sum_{i=1}^{M}f($$\{y_{i}(t),y_{j}(t)\})$
$+\frac{1}{N^{LF}}$ $\sum_{i=1}^{N^{LF}}f(\{y_{i}(t),x_{j}(t)\})\Bigg{]}\mathrm{d}t$
$+\sqrt{2D}\,\mathrm{d}W_{i},\;i=1,\dots,M.$ (56b)
Here, $x_{i}$ are positions of leaders and plastic followers, $y_{i}$ are positions of non-plastic followers, $N^{LF}$ and $M$ are their respective numbers, and $W_{i}$ is a standard Wiener process. Control input $u_{i}(t)=u(x_{i},t)$ is computed via spatial sampling, and $\lambda_{i}\in\{0,1\}$ indicates if agent $i$ is a leader ($\lambda=1$) or follower ($\lambda=0$). The label $\lambda_{i}$ is updated stochastically: leaders switch to plastic followers with rate $\kappa_{LF}(x_{i},t)$, plastic followers switch to leaders with rate $\kappa_{FL}(x_{i},t)$, where such space- and time-dependent rates are found factorizing $q$ in (17) as
$q(x,t)=\kappa_{FL}(x,t)\rho^{F}(x,t)-\kappa^{LF}(x,t)\rho^{L}(x,t).$ (57)
(a) $t=0$
(b) $t=t_{\mathrm{f}}$
(c) $t=t_{\mathrm{f}}$
Figure 3: Agent-based bimodal regulation. (a, b) Initial/final collective densities (solid black). In the inset, we show the discrete displacement of agents - leaders in blue, plastic followers in orange, non-plastic followers in purple (agents are plotted on concentric circles for visualization purposes). (c) Steady-state densities of the three populations (same color coding). (d) Time evolution of the KL divergence (top panel) and leaders’ and followers’ mass (bottom panel).
We consider a setup analogous to Sec. 4.1 with $N=1000$ agents (300 initial leaders, 300 initial plastic followers, 400 non-plastic followers), using Euler-Maruyama integration with $\Delta t=10^{-3}$ with $t_{\mathrm{f}}=15$. Control inputs $u$ and $q$ are computed via kernel density estimation from the agents’ positions (using the Matlab function circ_ksdensity with the msn method [43] and a width of 50 on a mesh of 600 points). Other parameters are set to $D=0.05$, $K=1$, $K_{FL}=2$, and $K_{LF}=1$.
In Fig. 3, we show a realization of (56). As one may expect, we register a performance degradation compared to the continuum case (see Fig. 1), due to the finite-size effect of the discretization. The steady-state leaders-to-followers ratio is approximately 1/2, consistent with Corollary 6. Due to the stochastic nature of (56), we repeat the same simulation scenario 50 times, recording an average KL divergence at steady-state of approximately 0.02 (±$\pm 0.01$), and an average leaders-to-followers mass ratio of 0.48 ($\pm$ 0.1).
6 Higher-Dimensional Extension
Next, we extend the theoretical framework to periodic domains in higher-dimensions, $\Omega=[-\pi,\pi]^{d}$ ($d=2,3$).
The model in (1) becomes
$\rho^{L}_{t}(\mathbf{x},t)+\nabla\cdot[\rho^{L}(\mathbf{x},t)$ $\mathbf{u}(\mathbf{x},t)+\rho^{L}(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)]=$
$=D\nabla^{2}\rho^{L}(\mathbf{x},t)+q(\mathbf{x},t),$ (58a)
$\rho^{F}_{t}(\mathbf{x},t)+\nabla\cdot[\rho^{F}(\mathbf{x},t)$ $(\mathbf{f}*\rho)(\mathbf{x},t)]=D\nabla^{2}\rho^{F}(\mathbf{x},t)$
$\qquad\qquad-q(\mathbf{x},t),$ (58b)
$\eta^{F}_{t}(\mathbf{x},t)+\nabla\cdot[\eta^{F}(\mathbf{x},t)$ $(\mathbf{f}*\rho)(\mathbf{x},t)]=D\nabla^{2}\eta^{F}(\mathbf{x},t),$ (58c)
where $\mathbf{f}$ is a $d$-dimensional periodic kernel and $\rho=\rho^{L}+\rho^{F}+\eta^{F}$. The system is complemented with periodic boundary conditions and initial conditions similar to (1).
6.1 Bio-inspired Control
Theorem 9.
Assume $\rho^{L}>0$ for any $x\in\Omega$ and $t\in\mathbb{R}_{0}$. Choosing
$\nabla\cdot\left[\rho^{L}(\mathbf{x},t)\mathbf{u}(\mathbf{x},t)\right]=-Ke(\mathbf{x},t)\\ -\nabla\cdot\left[\rho(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]-D\nabla^{2}\rho(\mathbf{x},t)$ (59)
where $K>0$ is a control gain, the error dynamics converges globally and exponentially to 0 pointwise in $\Omega$.
{pf}
The error dynamics obeys the higher-dimensional extension of (11). Choosing (59) brings the error dynamics in the form of (14), proving the claim. ∎
Remark 10.
To uniquely recover $\mathbf{u}$, a vector field, from the scalar relation (59), extra constraints need to be included, simlar to [44]. Specifically, we set $\mathbf{w}:=\rho^{L}\mathbf{u}$, and
$Y(\mathbf{x},t)=-Ke(\mathbf{x},t)-\nabla\cdot\left[\rho(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]\\ -D\nabla^{2}\rho(\mathbf{x},t),$ (60)
so that we can pose the problem
$\begin{cases}\nabla\cdot\mathbf{w}(\mathbf{x},t)=Y(\mathbf{x},t),\\ \nabla\times\mathbf{w}(\mathbf{x},t)=0,\end{cases}$ (61)
where we added a zero-curl condition to (59). Such a problem is analogous to the Poisson equation $\nabla^{2}\varphi=-Y$, where $\mathbf{w}=-\nabla\varphi$. The Poisson problem can be solved using Fourier series expansion, so that $\varphi$, and consequently $\mathbf{w}$ are recovered, from which one establishes $\mathbf{u}=\mathbf{w}/\rho^{L}$.
Analogously to the one-dimensional case, the resulting control velocity field $\mathbf{u}$ is well defined only if $\rho^{L}$ is strictly positive. Such a constraint can be ensured by appropriately choosing the reacting function $q$. In particular, extending (17) to higher dimensions, we establish
$q(\mathbf{x},t)=\frac{1}{2}\,\nabla\cdot\left[\rho^{L}(\mathbf{x},t)(\mathbf{x},t)\right]+\frac{1}{2}\,\nabla\cdot\left[\rho^{*}(\mathbf{x},t)(\mathbf{f}*\rho)(\mathbf{x},t)\right]\\ -\frac{D}{2}\nabla^{2}\rho^{*}(\mathbf{x},t)+g(\mathbf{x},t),$ (62)
where $\rho^{*}=\rho^{L}-\rho^{F}$ and $g$ is the mass action law in (19).
6.2 Stability Analysis
Theorem 11.
Assume the interaction kernel to be isotropic, that is,
$\int(f_{1}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{1}=\int(f_{2}*\psi)(x_{1},x_{2}x_{3})\,\mathrm{d}x_{2}=\\ =\int(f_{3}*\psi)(x_{1},x_{2},x_{3})\,\mathrm{d}x_{3},$ (63)
for any periodic $\psi$. Choosing $\mathbf{u}$ according to (59) (see Rem. 10) and $q$ as in (62) implies that $\bar{\rho}$ is a steady-state solution for the dynamics of $\rho$, with $\rho^{L},\rho^{F}>0$, and $\eta^{F}\geq 0$, if and only if the higher-dimensional extension of (20) holds.
{pf}
Under the additional assumption of isotropic interaction kernel, the proof follows the same steps of those in Theorem 4. The only difference is the computation of the steady-state solution of $\eta^{F}$ (see (27)). Setting $\eta^{F}_{t}=0$ and $\rho=\bar{\rho}$ in (58c), we obtain
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x})\right]$ $=D\nabla^{2}\bar{\eta}^{F}(\mathbf{x}),$ (64)
which is rewritten as
$\nabla\cdot\left[\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x})-D\nabla\bar{\eta}^{F}(\mathbf{x})\right]=0.$ (65)
Equation (65) is fulfilled if [^2] This condition becomes also necessary when $\bar{\eta}^{F}$, $\mathbf{f}$ and $\bar{\rho}$ satisfy ×$\nabla\times\left[\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}\right]=0$. This is equivalent to requiring that $\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}=-\nabla\Psi$, where $\Psi$ is a harmonic scalar potential. Under our isotropic kernel assumption, the solution we construct satisfies this condition, ensuring the uniqueness of the steady-state solution (see the following analysis and Remark 12).
$\nabla\bar{\eta}^{F}(\mathbf{x})=\frac{1}{D}\bar{\eta}^{F}(\mathbf{x})(\mathbf{f}*\bar{\rho})(\mathbf{x}).$ (66)
Equation (66) is a vectorial differential relation involving the partial derivatives of the scalar unknown $\bar{\eta}^{F}$, thus resulting in the ill-posed problem
$\begin{cases}\bar{\eta}^{F}_{x_{1}}(x_{1},x_{2})=\frac{1}{D}\bar{\eta}^{F}(x_{1},x_{2})(f_{1}*\bar{\rho})(x_{1},x_{2}),\\ \bar{\eta}^{F}_{x_{2}}(x_{1},x_{2})=\frac{1}{D}\bar{\eta}^{F}(x_{1},x_{2})(f_{2}*\bar{\rho})(x_{1},x_{2}).\end{cases}$ (67)
Here, without loss of generality, we set $d=2$, $\mathbf{f}=[f_{1},f_{2}]$, $\mathbf{x}=[x_{1},x_{2}]$ (the case $d=3$ is a trivial extension). We can now solve the two components of (67) separately, and check under which conditions they are equal.
By solving the first component of (67), we establish
$\bar{\eta}^{F}(x_{1},x_{2})=C_{1}(x_{2})\mathrm{exp}\left\{\frac{1}{D}\int(f_{1}*\bar{\rho})(x_{1},x_{2})\,\mathrm{d}x_{1}\right\}.$ (68)
where $C_{1}$ is a function of $x_{2}$ resulting from the spatial integration with respect to $x_{1}$. Similarly, if we integrate the second equation of (67), we get
$\bar{\eta}^{F}(x_{1},x_{2})=C_{2}(x_{1})\mathrm{exp}\left\{\frac{1}{D}\int(f_{2}*\bar{\rho})(x_{1},x_{2})\,\mathrm{d}x_{2}\right\}.$ (69)
where $C_{2}$ is a function of $x_{1}$ resulting from the spatial integration with respect to $x_{2}$.
For (68) and (69) to be equal, $C_{1}(x_{2})=C_{2}(x_{1})=C$, and the isotropic hypothesis (63) must hold. The value of $C$ can finally be chosen so that $\bar{\eta}^{F}$ integrates to $\Phi^{F}$ (note that the steady-state solution of (58c) is in the same form of its 1D counterpart - see (27)). The remainder of the proof follows that of Theorem 4.∎
Remark 12.
Many interaction kernels from the literature satisfy condition (63). Also, under this hypothesis, the steady-state solution for $\bar{\eta}^{F}$ is uniquely defined in higher dimensions, since $\bar{\eta}^{F}(\mathbf{f}*\bar{\rho})-D\nabla\bar{\eta}^{F}=-\nabla\Psi$, where $\Psi$ is a harmonic scalar potential.
Next, we extend Theorem 8 to higher dimensions.
(a) $t=t_{\mathrm{f}}$
(b) $t=t_{\mathrm{f}}$
(c) $t=t_{\mathrm{f}}$
Figure 4: 2D Mono-modal regulation. Final densities of (a) leaders’, (b) plastic followers’, and (c) non-plastic followers. (d) Time evolution of the KL divergence between $\rho$ and $\bar{\rho}$ (truncated to 10 time units for visualization purposes).
Theorem 13.
Under the conditions of Theorem 11 and if
$\sum_{i=1}^{d}\|\bar{\rho}_{x_{i}}(\cdot)\|_{2}\|f_{i}(\cdot)\|_{2}<2D,$ (70)
the error dynamics locally converges to 0 in $\mathcal{L}^{2}(\Omega)$.
{pf}
The proof follows the same structure of that of Theorem 8. Dropping time and space dependencies for simplicity, the time derivative of the Lyapunov functional in (44) can be rewritten as
$V_{t}=-2D\|\nabla e^{\eta}\|_{2}^{2}-\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\\ -2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x},$ (71)
using the divergence theorem, vectorial identities, and posing $\tilde{\mathbf{v}}=\nabla\cdot(\bar{\eta}^{F}\,\mathbf{f}*e^{0})$. The first term on the right-hand side can be bounded using Poincaré-Wirtinger inequality, and for the second one the following bound holds:
$\left|\int_{\Omega}\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\,\mathrm{d}\mathbf{x}\right|\leq\int_{\Omega}\left|\left(e^{\eta}\right)^{2}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right|\,\mathrm{d}\mathbf{x}=\\ =\left\|e^{\eta}e^{\eta}\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\right\|_{1}\leq\|e^{\eta}\|_{2}^{2}\|\nabla\cdot\left(\mathbf{f}*\bar{\rho}\right)\|_{\infty}\leq\\ \leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|(f_{i}*\bar{\rho}_{x_{i}})\|_{\infty}\leq\|e^{\eta}\|_{2}^{2}\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2},$ (72)
[…]
$\left|-2\mathrm{exp}(-Kt)\int_{\Omega}e^{\eta}\tilde{\mathbf{v}}\,\mathrm{d}\mathbf{x}\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{\mathbf{v}}\|_{1}\leq\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{\mathbf{v}}\|_{2}$ (73)
This leads us to the following bound on the time derivative of the Lyapunov functional
$V_{t}\leq\left(-2D+\sum_{i=1}^{d}\|f_{i}\|_{2}\|\bar{\rho}_{x_{i}}\|_{2}\right)V\\ +2\mathrm{exp}(-Kt)\|\tilde{\mathbf{v}}\|_{2}\sqrt{V}.\qed$ (74)
6.3 Numerical validation
We consider a 2D mono-modal regulation scenario with a bivariate von Mises distribution (zero means, unit concentration). Agents interact through a 2D periodic Morse kernel ($L_{a}=\pi$, $L_{r}=\pi/4$, $\alpha=3.2$) with $D=0.05$, $\Phi^{F}=0.2$, $M^{L}(0)=M^{F}(0)=0.4$, $K=1$, $K_{FL}=1$, $K_{LF}=2$. Results in Fig. 4 for $t_{\mathrm{f}}=100$ are qualitatively comparable to those of 1D simulations in Figs 1. Convergence of $\rho$ to $\bar{\rho}$ occurs in  5 time units while convergence of $\rho^{L}$, $\rho^{F}$ and $\eta^{F}$ to their steady-state profiles is slower. We obtain final masses $M^{L}(t_{\mathrm{f}})\approx 0.26$ and $M^{F}(t_{\mathrm{f}})\approx 0.53$, consistent with the predicted $1/2$ ratio of in Corollary 6.
7 Conclusions
We presented a novel bio-inspired leader-follower technique to control the spatial organization of large swarms of mobile agents. Upon formulating a density control problem, we derived conditions for the existence and local stability of desired solutions. Our strategy incorporates two crucial, yet previously overlooked, characteristics of natural systems. Namely, we set control objectives for the entire collective, including both leaders and followers, and we introduce behavioral plasticity to allow for tuning the leaders-to-follower mass ratio. We provide necessary and sufficient conditions for the existence of solutions to our control problem and assess their local stability, both for 1D and higher-dimensional settings. Our numerical findings offer compelling support to the mathematical derivations and to the effectiveness, robustness, and versatility of the proposed bio-inspired control strategy.
Our work does not come without limitations. Our analytical guarantees of convergence hold at the scale of PDEs, that is, in the limiting scenario of swarms of infinite size. The convergence from the continuum macroscopic PDE model to the discrete microscopic agent-based implementation should be assessed; potentially, this may be tackled through two-scale [45] and $\Gamma$-convergence [46]. Moreover, our theoretical results yield exact predictions only for steady-state leaders-to-followers mass ratio. Although numerical simulations suggest a monotonic trend in the leaders’ and plastic followers’ masses, we cannot argue that these trends are universal. As such, we may not exclude cases in which there is considerably higher role-switching during the system evolution than steady-state predictions. Exploring these transient dynamics may be valuable in the characterization of energetic costs of behavioral plasticity. Beyond addressing these limitations, we envision future work in which continuum descriptions of agents are not solely used for spatial organization, but to accomplish more complex tasks, such as self-assembly, collaborative manipulation, and object clustering [47].
{ack}
This work was supported in part by Italian Ministry of University and Research (MUR) performing the Activities of the Project PRIN 2022 MENTOR and in part by the NSF Grants CMMI-1932187 and EF-2222418.
Appendix A Non-plastic followers at steady-state
Theorem 14.
If $\rho(x,t)=\bar{\rho}(x)$, (1c) admits only the steady-state solution
$\bar{\eta}^{F}(x)=\frac{\Phi^{F}}{\int_{\mathcal{S}}h(x)\,\mathrm{d}x}h(x),$ (75)
with $h$ defined as in (21).
{pf}
Substituting $\eta^{F}_{t}=0$, $\eta^{F}(x,t)=\bar{\eta}^{F}(x)$, and $\rho(x,t)=\bar{\rho}(x)$ into (1c) leads us to (26). Integrating in space and isolating $\bar{\eta}^{F}_{x}$ at first member, we find
$\bar{\eta}^{F}_{x}(x)=\frac{\bar{\eta}^{F}(x)(f*\bar{\rho})(x)}{D}+A,$ (76)
where $A$ is an integration constant. The solution of (76) can be written as
$\bar{\eta}^{F}(x)=B\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(x)\,\mathrm{d}x\right]+\\ +A\,\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(x)\,\mathrm{d}x\right]\\ \times\int\mathrm{exp}\left[\frac{1}{D}\int(f*\bar{\rho})(y)\,\mathrm{d}y\right]\,\mathrm{d}x,$ (77)
where $B$ is an integration constant.
$\bar{\eta}^{F}$ must be periodic. The first term on the right-hand-side of (77) is positive and periodic (it is the exponential of a periodic function). Notice that $f*\bar{\rho}$ is periodic, as a result of the circular convolution, and it sums to 0 when integrated over $\mathcal{S}$ due to Fubini’s theorem for convolutions. Hence, the integral of $f*\bar{\rho}$ is itself periodic. The second term on the right-hand-side of (77), cannot be periodic unless $A=0$, since $\mathrm{exp}\left\{\frac{1}{D}\int(f*\bar{\rho})(y)\,\mathrm{d}y\right\}$ is periodic, but it cannot sum to 0, being an exponential. Thus, $A=0$. For (77) to integrate to the non-plastic followers’ mass $\Phi^{F}$, we must have $B=\Phi_{F}/\int_{\mathcal{S}}h(x)\,\mathrm{d}x$, yielding the claim.∎
Remark 15.
(75) is positive by construction, as $h$ is an exponential (see (21)) and $\frac{\Phi_{F}}{\int_{\mathcal{S}}h(x)\,\mathrm{d}x}>0$ by construction.